OpenStudy (anonymous):

Find the value of x so that the line passing through (10, 5) and (x, 9) has a slope of -2.

OpenStudy (anonymous):

To find slope: \[\frac{ y _{2} - y _{1} }{ x _{2} - x _{1} }\] Plug in your given values: \[\frac{ 9 - 5 }{ x - 10 }\] Simplify: \[\frac{ 4 }{ x - 10 }\] And you can solve it for x from there.

OpenStudy (anonymous):

X=40?